Database last updated on Sun, 01 Feb 2026 20:10:55 +0000 by localhost

Instrument constant changes for the week before 2024030 (5 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 13 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
085202403021026.4

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
0852024030479.9

Non-critical warnings from b-files' comments for days between 2024030 and 2024030 (0 sec, 4 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
user:Brewer Restarted199B03024.085TORON#08508:45:1900
fr:Micrometer #1 reference found at step 9142 (+CAL)= 10166 Now set to o3 offset = 7832199B03024.085TORON#08513:38:421016610009000
hg:DEL key pressed299B03024.085TORON#08515:58:501-1
ch:Command file ch not found199B03024.085TORON#08520:04:46found1-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssluv
030TORONTO#085Going to menu,inside,menu,Must do HG now,Waiting until ZA =,Waiting until ZA = 170 now 148.1,Waiting until ZA = 170 now 149.3,Waiting until ZA = 170 now 150.3,Waiting until ZA = 170 now 151.2,Waiting until ZA = 170 now 152,Waiting until ZA = 170 now 152.7,Waiting until ZA = 170 now 153.2,Waiting until ZA = 170 now 153.5,Waiting until ZA = 170 now 153.72210078221
(1 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B03024.0852#0852829